OpenStudy (anonymous):

how do you convert 11.5 miles to centimeters? show the process please!

OpenStudy (phi):

do you have a "conversion factor" for miles to some other unit (not necessarily cm) ?

OpenStudy (anonymous):

i not sure what that means?

OpenStudy (phi):

Have you learned how many miles in a km ? or how many km in a mile ?

OpenStudy (anonymous):

aren't there a thousand km in a mile

OpenStudy (phi):

no. there is a thousand meters in a kilometer.

OpenStudy (phi):

But let's use 1 mile = 1.609 kilometers to convert 11.5 miles to km one way to remember is to say: I want to "cancel" miles and replace with km in algebra we would do \[ mile \cdot \frac{km}{mile} \\ =\cancel{mile} \cdot \frac{km}{\cancel{mile} }= km \]

OpenStudy (phi):

we start with 1 mile = 1.609 kilometers and write it as a ratio: \[ \frac{ 1.609\ km}{1 \ mile} \] notice we wrote it with km on top and mile in the bottom. now multiply \[ 11.5\ mile \cdot \frac{ 1.609\ km}{1 \ mile} \]

OpenStudy (phi):

we have mile in the top and bottom, so we cancel them we are left with \[ 11.5 \cdot 1.609 \ km\] use a calculator to multiply the numbers. we now have the number of km

OpenStudy (phi):

how many km do we have ?

OpenStudy (anonymous):

i don't need km i need it in cm sorry if i wasn't clear about that

OpenStudy (phi):

Yes, but once we have it in km, we can change to meters and then into cm

OpenStudy (anonymous):

how do i do that

OpenStudy (phi):

first, how many km do we have ?

OpenStudy (anonymous):

18.5 km

OpenStudy (phi):

we probably want more digits. In fact, let's use all of them 18.5035 km

OpenStudy (phi):

I told you up above, how many meters in 1 km ?

OpenStudy (anonymous):

1000

OpenStudy (phi):

yes. Because we are converting between units, get used to writing it *all out* 1000 m = 1 km 1000 meters is 1 kilometer as a step to changing to cm, we will change to m (and then to cm) we have 1000 m = 1 km First, we write this as a ratio we have 2 choices \[ \frac{ 1000 \ m}{1 \ km} \\ \text{ or } \frac{1 \ km}{1000 \ m} \] which ratio should we use to change 18.5035 km to m ?

OpenStudy (phi):

remember we want m up top and km in the bottom (so it will cancel the other km)